# Where is the Truth? The Risk of Getting Confounded in a Continual World

### v2

@@ C.2 Training Converges on Confounded Data with Worse Results on Unconfounded Data
Figure 5: Learning curves for the cumulative training setup. Recall that in the cumulative setting the dataset being trained on grows over time, i.e., all datasets from previous tasks are concatenated with the currently observed one. This figure compares these curves to the test accuracy on the unconfounded dataset (denoted by Unconf.), task-specific confounded datasets (denoted by Conf.), and average accuracy calculated over accumulated confounded tasks (denoted by Conf. Avg.) at the end of the training. Average accuracy after $T$ tasks is given as: $A_{T}=\frac{1}{T}\sum_{t}^{T}a_{t}$. Observe that in the strict case, performance on the unconfounded data does increase during training, but the final performance is still poor. We expect that this can be explained by the further observation of the divergence between train and validation curves in the strict case.
Recall that, in the strict setting, we saw the models and the NN in particular achieve a high accuracy on the unconfounded dataset in the joint setting, while training on the same data in a cumulative manner resulted in an accuracy more than 20 points worse. This result is surprising since both setups are trained on the exact same datasets. Hence, one might wonder whether the model actually converged to a good accuracy during training in the cumulative setup. As we can see in Figure 5, both models indeed converge at close to perfect accuracy. For example, the NN ends at 100% train accuracy after task 1 and 99.5% accuracy after task 2. Training on the full accumulated data in task 3 peaks at 99.96% and ends at 99.68% accuracy. Despite this, performance on the unconfounded dataset, while increasing with the number of tasks, remains significantly worse than the train accuracy. In the main paper, we showed that the easiest rule that can be learned to solve all confounded tasks jointly is the ground truth rule. If that is true, why does the NN achieve a near-perfect train accuracy after training on all tasks but fail on the unconfounded data where the ground truth also applies? By inspecting the validation and test values of the confounded datasets, we find that, despite coming from the same distribution of images, accuracies on these are considerably worse than for the train images. It turns out that the cumulative setting has resulted in both models overfitting far more severely to the train distribution. It thus underperforms on both confounded data from the same distribution and unconfounded data due to a general failure to fully generalize to new images. We expect that more investigation of this overfitting will further illuminate the problem of insidious continual confounding.
@@ C.3 Models Mostly Fail on Positive Images in disjoint Settings
Figure 6: Accuracies for positive and negative images on disjoint. Test accuracies on both the NN and NeSy. We can see that the prediction error is due to the positive images which are incorrectly classified as negative. This is consistent with our hypothesis of the model mostly basing its predictions on the presence of the confounding features, none of which appear in any negative images of the confounded tasks. On the unconfounded data, the high performances on positive images after tasks 2 and 3 indicate that a disjunction of confounding features is learned since this is satisfied in most positive and only few negative images. (Best viewed in color)
@@
We employ a new visualization, extending the purely tabular results presented before. Here, we measure the accuracy on the confounded test dataset across all the tasks after training the model on each task. To investigate whether and how the model gets confounded, we also measure the accuracy on the unconfounded dataset. These results are illustrated via an inter-task confusion matrix in Figure 6. Each row shows the results for a continual learning setup on both NN (left) and NeSy (right). Instead of the average accuracies already discussed in the main paper, we here plot the results on the positive and negative sets of images. We additionally color the fields corresponding to their accuracy, going from incorrect predictions (0%; red) over random predictions (50%; blue) to correct predictions (100%, white).
@@
The results of GDUMB on NN set themselves apart from the other methods. We will quickly cover the other methods and then return to GDUMB. With the exception of GDUMB, predictions on the negative set across all tasks and methods are mostly or even fully correct, and the classification errors are almost exclusively concentrated on the positive set. In other words, the models seem to predict the negative class too often, hence failing on the positive set. This can be explained by the model predicting the positive class whenever the confounding feature is present (it is never present on negative images of confounded tasks). The model also performs well on all previous tasks in FROMP, ER, DER, BGS, and the cumulative setting, indicating that some combination of confounders is learned.
@@
We can get some more insights into the model behavior by relating it to the performance on the unconfounded dataset. For all methods, performance on both positive and negative images is around 50% after the first task. Here, the model appears to make predictions based on the first confounding feature, blue, as in the unconfounded dataset, this roughly matches the proportion of images with and without a blue object. However, almost all images are predicted to be positive for the two following tasks. Our results on the memory-based methods FROMP, ER, DER, BGS, GDUMB, and the cumulative setup, which show good performance on all current and previously seen tasks, lead us to hypothesize that the disjunction of all confounding features is learned. In any image of the unconfounded dataset it is very likely that at least one confounding feature is present, hence resulting in predictions for the positive class. For the naive setup and EWC, the higher accuracies on the positive images can also partly be explained by the data generation process: for both metal and large, it is more likely for an unconfounded image to contain the respective features than not to contain them.
@@
Overall, results on the positive and negative images show that the negative set of confounded tasks can be consistently predicted correctly. This is entirely plausible and perhaps even expected as the negative set used in all tasks follows shared rules for generation. In turn, errors are introduced by incorrect predictions on positive images. The accuracies on the unconfounded dataset conform with our hypothesis of the models learning a disjunction of confounders, a rule that is mostly satisfied on positive unconfounded images.
@@ C.4 Models Base Predictions on Task-Specific Confounders in strict Settings
Figure 7: Accuracies for positive and negative images on strict in resemblance to Figure 6. After task 1, performance on positive and negative images is relatively balanced. Here, predictions based on the confounding feature perform well on neither positive or negative images of other tasks since these feature appear there randomly. For tasks 2 and 3, the respective confounding features are more likely to be present, so the model tends to classify images as “positive”. This is due to the first confounding feature (blue) having more alternative values (there are eight colors) than the confounding features metal and large, which both have only one alternative property value. (Best viewed in color)
@@
For the strict dataset, our evaluation in the main paper demonstrated the problematic manifestation of insidious continual confounding, i.e. a model’s performance when trained on all confounded task jointly is considerably higher than when trained on the same data in a cumulative, continual setup. In the latter, it largely fails.
@@
Figure 7 shows the individually decomposed results for the strict dataset variant, analogous to the analysis of the previous subsection. Again, the model appears to have learned the first confounder after the first task. However, due to this confounding feature being allowed to be present in other tasks, this now leads to more balanced predictions on future tasks. After training on tasks 2 and 3, most predictions fail on the negative instead of the positive sets of images. This is again consistent with our hypothesis of the model learning the task-specific confounder. To be precise, both the confounders of tasks 2 and 3 describe a property, which, in our dataset, only has two different possible values (metal or rubber; small or large). Since these properties are determined randomly if not specified, there is only a small probability of that feature not being present [^6] Each of the four objects in an unconfounded image has a chance of containing the material metal. The probability of such an image not containing any metal object is, therefore, quite low. The same idea applies analogously to large (on positive images, the small sphere part of the ground truth rule also affects the distribution of sizes).. Therefore, applying the rule based on the task-specific confounding feature mostly classifies any such image as positive, no matter the class, thus adding up to around 50% on average.
@@
Despite cumulative training, which performs better on old tasks of both classes but still fails on a large number of unconfounded images, there are two methods with results different from the others. FROMP has worse results on positive but better results on negative images. Here, old task accuracy is improved at the cost of current task performance. While this indicates that something different from simply the current task confounder is learned, the unconfounded accuracy remains to be not much higher than 50%. The other outlier method, GDUMB, performs better on negative images but much worse on positive ones. Since just like with the cumulative training, the ratio of samples from each class is the same, it does not prioritize the current task confounder. However, it still fails to make correct predictions on many images of the confounded tasks and also gives bad predictions on the unconfounded dataset.
@@
Appendix D Limitations and Impact
@@
Limitations. ConCon is the first dataset designed to investigate how ML and CL methods behave in the case of continual confounding. Nevertheless, there are certain possible research directions on continual confounding variations where an extension of the current ConCon could be beneficial. While our code framework for generating datasets can be changed to allow for different kinds of continual confounding, the current implementation only aims to generate problems where confounders only appear on a single object. Multi-object confounders are not implemented as of now. Other types of confounders, such as relations between objects or positioning (e.g., there is a sphere left to a red object, there is a cube on the far left), are also not considered yet. Generally, any CLEVR-based dataset will have limitations due to its synthetic nature. Investigating continually confounded real-world datasets is thus left for future work. In addition, ConCon only considers confounding in a visual domain. Other modalities, such as text or time series, are not currently in the focus of ConCon.
Impact Statement. Our paper challenges the conventional view of cumulative learning as an upper bound when evaluating continual learning methods, thereby showing that such confounders have a detrimental effect on model performance even in what is conventionally considered a best-case scenario. The issue of confounders in machine learning represents a significant challenge. Making predictions based on such can not only lead to incorrect behavior but may also lead to unfair or discriminatory decisions [41]. In our specific case of ConCon, the dataset consists of synthetically generated geometric shapes and thus does not contain any personal or otherwise private information. At present, we do not envision any scenario in which it could be used with malicious intent. As such, we hope that this work and investigation with the ConCon dataset will help improve continual learning algorithms in the future, especially with respect to how memory is being used.
Author Statement of Responsibility
The authors accept their responsibility in case of violation of rights related to the dataset provided concerning copyrighted material or data privacy. We confirm that we have, to the best of our knowledge, obtained all necessary permissions and in good faith complied with all relevant legal and ethical requirements in the preparation and submission of this dataset.
Appendix E ConCon Datasheet
@@
The creators of the dataset are: Florian Peter Busch1,2, Roshni Ramanna Kamath1,2, Rupert Mitchell1,2, Wolfgang Stammer1,2, Kristian Kersting1,2,3,4 and Martin Mundt1,2. The associated affiliations are: 1Department of Computer Science, TU Darmstadt, Darmstadt, Germany, 2Hessian Center for AI (hessian.AI), Darmstadt, Germany, 3German Research Center for Artificial Intelligence (DFKI), Darmstadt, Germany and 4Centre for Cognitive Science, TU Darmstadt, Darmstadt, Germany.
@@
The dataset is self-contained, however it does build on prior work. It is built on the CLEVR framework (Johnson et al. 2017, CVPR), which has an open source license that allows us to make these modifications. Data generation also relies on the publicly accessible and free Blender software.
@@
The source code used to generate the datasets is available at:
https://github.com/ml-research/concon.

### v3

@@ C.2 Training Converges on Confounded Data with Worse Results on Unconfounded Data
Figure 5: Learning Curves for the Cumulative Training Setup. Recall that in the cumulative setting, the dataset being trained on grows over time, i.e., all datasets from previous tasks are concatenated with the currently observed one. This figure compares these curves to the test accuracy on the unconfounded dataset (denoted by Unconf.), task-specific confounded datasets (denoted by Conf.), and average accuracy calculated over accumulated confounded tasks (denoted by Conf. Avg.) at the end of the training. Average accuracy after $T$ tasks is given as: $A_{T}=\frac{1}{T}\sum_{t}^{T}a_{t}$. Observe that in the strict case, performance on the unconfounded data does increase during training, but the final performance is still poor. We expect that this can be explained by the further observation of the divergence between train and validation curves in the strict case.
Recall that, in the strict setting, we saw the models and the NN in particular achieve a high accuracy on the unconfounded dataset in the joint setting, while training on the same data in a cumulative manner resulted in an accuracy more than 20 points worse. This result is surprising since both setups are trained on the exact same datasets. Hence, one might wonder whether the model actually converged to a good accuracy during training in the cumulative setup. As we can see in Figure 5, both models indeed converge at close to perfect accuracy. For example, the NN ends at 100% train accuracy after task 1 and 99.5% accuracy after task 2. Training on the full accumulated data in task 3 peaks at 99.96% and ends at 99.68% accuracy. Despite this, performance on the unconfounded dataset, while increasing with the number of tasks, remains significantly worse than the train accuracy. In the main paper, we showed that the easiest rule that can be learned to solve all confounded tasks jointly is the ground truth rule. If that is true, why does the NN achieve a near-perfect train accuracy after training on all tasks but fail on the unconfounded data where the ground truth also applies? By inspecting the validation and test values of the confounded datasets, we find that, despite coming from the same distribution of images, the accuracies on these are considerably worse than for the train images. It turns out that the cumulative setting has resulted in both models overfitting far more severely to the train distribution. It thus underperforms on both confounded data from the same distribution and unconfounded data due to a general failure to fully generalize to new images. We expect that more investigation of this overfitting will further illuminate the problem of insidious continual confounding.
@@ C.3 Models Mostly Fail on Positive Images in disjoint Settings
Figure 6: Accuracies for Positive and Negative Images on disjoint. Test accuracies on both the NN and NeSy for Naive Sequential, Shuffled, and Cumulative. We can see that the prediction error is a result of positive images being incorrectly classified as negative. This is consistent with our hypothesis that the model mostly bases its predictions on the presence of the confounding features, none of which appear in any negative images of the confounded tasks. On the unconfounded data, the high performances on positive images after tasks 2 and 3 indicate that a disjunction of confounding features is learned since this is satisfied in most positive and only a few negative images. Further results on the other continual learning methods can be found in Figure 7. (Best viewed in color)
Figure 7: Accuracies for Positive and Negative Images on disjoint. Test accuracies on both the NN and NeSy for different continual learning methods. As for the baseline methods, we see that the model failures on confounded tasks are due to incorrectly predicted positive images, while too many unconfounded images are incorrectly predicted as belonging to the positive class. (Best viewed in color)
@@
We employ a new visualization, extending the purely tabular results presented before. Here, we measure the accuracy on the confounded test dataset across all the tasks after training the model on each task. To investigate whether and how the model gets confounded, we also measure the accuracy on the unconfounded dataset. These results are illustrated via an inter-task confusion matrix in Figures 6 and 7. Each row shows the results for a continual learning setup on both NN (left) and NeSy (right). Instead of the average accuracies already discussed in the main paper, we here plot the results on the positive and negative sets of images. We additionally color the fields corresponding to their accuracy, going from incorrect predictions (0%; red) over random predictions (50%; blue) to correct predictions (100%, white).
First, consider the shuffled approach. Here, tasks are not aligned with confounders, which is why performance on all tasks is very similar. Our hypothesis of the model learning a disjunction of confounders fits the results, as all tasks are predicted correctly but most negative images of the unconfounded datasets are incorrectly classified as positive. In the following, we go into further detail on the continual learning setups that are affected by confounders to further explain the models’ behavior there.
@@
The results of GDUMB on NN set themselves apart from the other methods. We will quickly cover the other methods and then return to GDUMB. With the exception of GDUMB, predictions on the negative set across all tasks and methods are mostly or even fully correct, and the classification errors are almost exclusively concentrated on the positive set. In other words, the models seem to predict the negative class too often, hence failing on the positive set. This can be explained by the model predicting the positive class whenever the confounding feature is present (it is never present on negative images of confounded tasks). The model also performs well on all previous tasks in FROMP, ER, DER, BGS, PNNs, and the cumulative setting, indicating that some combination of confounders is learned.
@@
We can get some more insights into the model behavior by relating it to the performance on the unconfounded dataset. For all methods, performance on both positive and negative images is around 50% after the first task. Here, the model appears to make predictions based on the first confounding feature, blue, as in the unconfounded dataset, this roughly matches the proportion of images with and without a blue object. However, almost all images are predicted to be positive for the two following tasks. Our results on the memory-based methods FROMP, ER, DER, BGS, GDUMB, as well as PNNS and the cumulative setup, which show good performance on all current and previously seen tasks, lead us to hypothesize that the disjunction of all confounding features is learned. In any image of the unconfounded dataset, it is very likely that at least one confounding feature is present, hence resulting in predictions for the positive class. For the naive setup and EWC, the higher accuracies on the positive images can also partly be explained by the data generation process: for both metal and large, it is more likely for an unconfounded image to contain the respective features than not to contain them.
@@
Overall, results on the positive and negative images show that the negative set of confounded tasks can be consistently predicted correctly. This is entirely plausible and perhaps even expected, as the negative set used in all tasks follows shared rules for generation. In turn, errors are introduced by incorrect predictions on positive images. The accuracies on the unconfounded dataset conform with our hypothesis of the models learning a disjunction of confounders, a rule that is mostly satisfied on positive unconfounded images.
@@ C.4 Models Base Predictions on Task-Specific Confounders in strict Settings
Figure 8: Accuracies for Positive and Negative Images on strict. Test accuracies on both the NN and NeSy for Naive Sequential, Shuffled, and Cumulative. For naive sequential, performance on positive and negative images is relatively balanced after task 1. For tasks 2 and 3, the respective confounding features are more likely to be present, so the model tends to classify images as “positive” in the naive sequential approach. This is due to the first confounding feature (blue) having more alternative values (there are eight colors) than the confounding features metal and large, which both have only one alternative property value. In the shuffled setup for NN, the accuracies are similar across tasks, with only the unconfounded accuracy improving with additional data. Cumulative also improves with additional data, but stays behind shuffled in terms of accuracy. (Best viewed in color)
Figure 9: Accuracies for Positive and Negative Images on strict. Test accuracies on both the NN and NeSy for different continual learning methods. Similarly to naive sequential, performance on positive and negative images is mostly balanced after task 1. In the second and third tasks, the models tend to classify images as “positive” more often, as confounding features are more likely to be present than not. (Best viewed in color)
@@
For the strict dataset, our evaluation in the main paper demonstrated the problematic manifestation of insidious continual confounding, i.e. a model’s performance when trained on all confounded tasks jointly is considerably higher than when trained on the same data in a cumulative, continual setup. In the latter, it largely fails.
@@
Figures 8 and 9 show the individually decomposed results for the strict dataset variant, analogous to the analysis of the previous subsection. As for the disjoint case, accuracies in the shuffled case are mostly unchanged throughout tasks because the tasks are not aligned with the confounding features. We can, however, see that additional data helps the model learn the ground truth rule, as accuracies on both positive and negative images go up in tasks 2 and 3. Considering the continual learning setups, we again see that the models appear to have learned the first confounder after the first task. However, due to this confounding feature being allowed to be present in other tasks, this now leads to more balanced predictions on future tasks. After training on tasks 2 and 3, most predictions fail on the negative instead of the positive sets of images. This is again consistent with our hypothesis of the model learning the task-specific confounder. To be precise, both the confounders of tasks 2 and 3 describe a property, which, in our dataset, only has two different possible values (metal or rubber; small or large). Since these properties are determined randomly if not specified, there is only a small probability of that feature not being present. [^8] Each of the four objects in an unconfounded image has a chance of containing the material metal. The probability of such an image not containing any metal object is, therefore, quite low. The same idea applies analogously to large (on positive images, the small sphere part of the ground truth rule also affects the distribution of sizes). Therefore, applying the rule based on the task-specific confounding feature mostly classifies any such image as positive, no matter the class, thus adding up to around 50% on average.
@@
While the results for the CL methods look mostly similar, there are three methods with results different from the others. FROMP has worse results on positive but better results on negative images. Here, old task accuracy is improved at the cost of current task performance. While this indicates that something different from simply the current task confounder is learned, the unconfounded accuracy remains to be not much higher than 50%. The second outlier method, GDUMB, performs better on negative images but much worse on positive ones. Since, just like with the cumulative training, the ratio of samples from each class is the same, it does not prioritize the current task confounder. However, it still fails to make correct predictions on many images of the confounded tasks and also gives bad predictions on the unconfounded dataset. Lastly, PNNs mostly fall in line with the other methods, but the task information allows them to correctly predict positive and negative images of all previous tasks. On the other hand, this still does not help with making correct predictions on the unconfounded data, where PNNs do not perform better than the other methods.
@@
Appendix D Experiment Using Generative Models
Table 6: Test Accuracies for Generative Replay on the ConCon Dataset. While accuracies on the current tasks are high, we observe catastrophic forgetting, resulting in low accuracies on old tasks. Accuracy on the unconfounded dataset is around $50\%$, showing that the model does not successfully learn the ground truth rule.
$T_{1}$ $T_{2}$ $T_{3}$ $T_{1}@T_{3}$ $T_{2}@T_{3}$ UnConf
disjoint 100.0$\pm$0.0 88.78$\pm$14.29 99.28$\pm$1.02 53.2$\pm$6.30 59.33$\pm$15.02 47.73$\pm$1.45
strict 99.96$\pm$0.08 98.54$\pm$0.51 99.58$\pm$0.46 48.84$\pm$0.48 48.61$\pm$0.56 49.26$\pm$0.88
In addition to the NN and NeSy models, we here want to investigate whether another type of model based on generative models is more successful in overcoming continual confounding. To this end, we use a Variational Autoencoder (VAE) architecture based on ResNet-18 following van de Ven et al. (2020). The VAE framework allows the model to generate synthetic samples of the ConCon dataset from previous tasks by sampling from the learned latent space, providing a mechanism for replay without storing actual historical data. This experiment uses the same hyperparameters as the experiment in the main body (see Appendix B). Table 6 shows the experimental results on both the disjoint and strict variants of our ConCon dataset. The model achieves high accuracies on the current tasks across both disjoint and strict settings, but fails at correctly predicting images from previous tasks, showcasing catastrophic forgetting. More critically (but, given the bad performance on previous tasks, also unsurprisingly), the model fails completely on the unconfounded dataset, indicating, like other continual learning methods, that it learns the task-specific confounders and not the ground truth rule.
Appendix E Expanded ImageNet Experiments
Following up on our motivational experiment using ImageNet classes at the start of the main body, we now conduct a larger experiment using ImageNet classes to further investigate whether we again observe a gap between joint and cumulative training. To this end, we consider two categories: four-legged animals and non-four-legged animals. We define three tasks where each task contains data from eight ImageNet classes, half of which are four-legged animals. Additionally, in each task, this group of four ImageNet classes shares a similarity with respect to color. For example, the two resulting classes in task 1 contain either four-legged animals where the predominant color is white (e.g., arctic fox, polar bear), while the second class contains four-legged animals where the predominant color is blue (e.g., jellyfish, blue jay). Therefore, when training the model in a continual manner (cumulative training), it is likely that the model will start by learning a simpler rule in task one, by, for example, making predictions mostly based on color, the “four-legged” feature, or even other features that we did not consider. Following our insights from the experiments of the main body, we hypothesize that such behavior will lead to the model having more trouble with correctly making predictions on all data compared to the joint setting, where it processes all the data at the same time and, hence, would not favor rules as much that are only consistent with a subset of all classes.
The exact description of the dataset and experimental results is shown in Table 7. We use a ResNet-18 model and average results across 5 random seeds (other hyperparameters are the same as for most other experiments, see Appendix B). The table shows the accuracies after the model has been trained on all 3 tasks. From joint training to cumulative training, we observe an average decrease in accuracy of 5.58 percentage points, a substantial degradation in performance. While the setup is slightly different from our continual confounding experiments in the main paper, it does showcase that the continual learning setting tends to result in a less accurate model than the joint setup.
Table 7: Cumulative and Joint Training on ImageNet Data. At the top, we show what ImageNet classes the dataset consists of. Each task includes a pair of classes defined by “4-leggedness” and dominant color, resulting in six classes overall. Class examples are from ImageNet. At the bottom, we report cumulative and joint training accuracies after training on all tasks. On average, joint training has an accuracy of 5.58 points higher than cumulative training.
Task Class Description ImageNet Classes
$T_{1}$ $C_{0}$ Four-legged, white color arctic fox, polar bear, white wolf, samoyed dog
$C_{1}$ Not four-legged, blue color jellyfish, blue jay, blue shark, tench
$T_{2}$ $C_{2}$ Four-legged, green color cheetah, deer, lion, leopard
$C_{3}$ Not four-legged, white color snowmobile, sailboat, snowplough, rattlesnake
$T_{3}$ $C_{4}$ Four-legged, blue color hippo, crocodile, water buffalo, beaver
$C_{5}$ Not four-legged, other colors eagle, spider, lifeboat, mushroom
Accuracy After Training on All Tasks
Training Setting $T_{1}$ $T_{2}$ $T_{3}$
Cumulative 76.30$\pm$0.45 78.50$\pm$2.69 84.40$\pm$0.86
Joint 83.45$\pm$2.46 86.25$\pm$1.24 86.25$\pm$1.15
Appendix F Limitations
@@
ConCon is the first dataset designed to investigate how ML and CL methods behave in the case of continual confounding. Nevertheless, there are certain possible research directions on continual confounding variations where an extension of the current ConCon could be beneficial. While our code framework for generating datasets can be changed to allow for different kinds of continual confounding, the current implementation only aims to generate problems where confounders only appear on a single object. Multi-object confounders are not implemented as of now. Other types of confounders, such as relations between objects or positioning (e.g., there is a sphere left to a red object, there is a cube on the far left), are also not considered yet. Generally, any CLEVR-based dataset will have limitations due to its synthetic nature. Investigating continually confounded real-world datasets thoroughly is thus left for future work.
Appendix G ConCon Datasheet
@@
The creators of the dataset are: Florian Peter Busch1,2, Roshni Kamath1,2, Rupert Mitchell1,2, Wolfgang Stammer1,2, Kristian Kersting1,2,3,4 and Martin Mundt5. The associated affiliations are: 1Department of Computer Science, TU Darmstadt, Darmstadt, Germany, 2Hessian Center for AI (hessian.AI), Darmstadt, Germany, 3German Research Center for Artificial Intelligence (DFKI), Darmstadt, Germany and 4Centre for Cognitive Science, TU Darmstadt, Darmstadt, Germany. 5University of Bremen, Germany.
@@
The dataset is self-contained, however it does build on prior work. It is built on the CLEVR framework (Johnson et al., 2017), which has an open source license that allows us to make these modifications. Data generation also relies on the publicly accessible and free Blender software.
@@
The source code used to generate the datasets is available at: https://github.com/ml-research/concon.